OpenStudy (anonymous):

Leonardo will pay $484 to three different credit cards: Expert Card, Passport, and US Speedy. He will pay four times as much to Expert Card as US Speedy since it has a higher balance. He also expects the Passport payment to be only 10% of the total of the other two payments. How much will Leonardo pay towards the Passport credit card balance?

OpenStudy (anonymous):

Love this question :)

OpenStudy (anonymous):

First, let set up some variables... US Speedy = U Expert = 4U Passport = .10(U+4U) = .10(5U) = .50U do you see where that all comes from in your word problems....

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great... Now, 484 = U + 4U +.50U solve for U - what do you get?

OpenStudy (anonymous):

all u do is add the Us up right

OpenStudy (anonymous):

yes add up U's; but then you must divide to get the term U with NO coefficien

OpenStudy (anonymous):

ya so i got 88 right

OpenStudy (anonymous):

absolutely correct....so lets go back to our variable set ups so speedy =88 what is express = 4u = what is passport .5U =

OpenStudy (anonymous):

so what do i do divide 484 by 4? and then by .5

OpenStudy (anonymous):

no wait thats not right

OpenStudy (anonymous):

4U = 4(88) = ????

OpenStudy (anonymous):

352

OpenStudy (anonymous):

you are now just replacing the variable U by the number you calculated of 88

OpenStudy (anonymous):

ooooo ok

OpenStudy (anonymous):

so its 44?

OpenStudy (anonymous):

right and right..... Great job!!!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

uw